OpenStudy (anonymous):

1/2(6-y)=y? one half that is.

OpenStudy (anonymous):

So do you mean\[\frac{1}{2}(6-y)=y\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you know the distributive property?

OpenStudy (anonymous):

Yes, I was just confused on what you do after you distribute.

OpenStudy (anonymous):

What did you get after you distribute the 1/2?

OpenStudy (anonymous):

3-1/2y=y

OpenStudy (anonymous):

Can you get the y alone on one side of the = ?

OpenStudy (anonymous):

So, you would subtract y from -1/2y?

OpenStudy (anonymous):

I think you mean subtract y from both sides (leaving 0 on the right). You could do that. What do you get when you do that?

OpenStudy (anonymous):

3-1/2=5/2

OpenStudy (anonymous):

Where did the y go?

OpenStudy (anonymous):

Never mind, I explained it wrong. Can we start over?

OpenStudy (anonymous):

From 3-1/2y=y

OpenStudy (anonymous):

If you go back to 3-(1/2)y = y, there's a -(1/2)y on the left that wants to be on the right. So try adding (1/2)y to both sides.

OpenStudy (anonymous):

I got the answer. Thank you!

OpenStudy (anonymous):

Post what you got so I can see.

OpenStudy (anonymous):

2=y

OpenStudy (anonymous):

Yep. 3-(1/2)y = y 3 = (3/2)y y = 3/(3/2) = 3*(2/3) = 2

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

OR, 6-y=2y (using cross multiplication, avoids working with fractions) 6=3y y=6/3=2